# Numerical modeling of two magnetized counter-propagating weakly collisional plasma flows in arch configuration

### v1

@@ Abstract
plasma hybrid simulation; particle-in-cell code; plasma arc; plasma discharge; magnetized plasma; weibel instability; magnetic field reconnection
@@ I Introduction
In astrophysics, the interaction of plasma flows with magnetic structures play a major role. In particular, in the atmospheres of stars and the magnetospheres of planets, magnetic field configurations in the form of an arch, usually filled with plasma, are widespread. Such magnetized plasma arch structures have been observed on the Sun Masuda et al. (1994) and in the Earth’s magnetosphere Liu (2020). It has been shown that they play a significant role in solar flares Xia et al. (2020), as well as in various explosive processes accompanied by the reconnection of magnetic lines Yamada et al. (2010).
@@
Despite the rich possibilities for performing various diagnostics in a laboratory experiment, many processes occurring in it have too small scales in time and space to be distinguished by modern methods. In this regard, numerical modeling is widely used to plan experiments and explain their results. As an example, we can cite a recent work on the numerical modeling of the interaction process of a plasma flow incident on a magnetic arch Wang et al. (2024) in a configuration similar to the experiment on our setup Viktorov et al. (2019a, b). At present, work is being carried out on the setup to study the collision process of two counter-streaming flows injected from the bases of a magnetic arch along magnetic lines. The present work is aimed at the numerical modeling of this process and the identification of the main interaction modes and the physical processes accompanying them. The main attention is paid to the emerging structures of plasma density, directly observed by its optical glow in the experiment, as well as to the processes of development of various instabilities and, in particular, the Weibel instability Weibel (1959).
@@ III Numerical methods
Numerical modeling of weakly collisional highly non-equilibrium plasma generally requires consistent consideration of kinetic effects, which can be ensured by numerical modeling of Vlasov—Maxwell system of equations Bertrand et al. (2019):
@@
In this regard, within the framework of this study, it was decided to conduct a full-scale modeling using a hybrid method, in which only ions are described kinetically, and electrons are represented as a massless neutralizing liquid. To partially account for kinetic effects, however, electrons were described in a so-called 10-moment approximation taking into account the evolution of their pressure tensor. We also neglect displacement current $\partial\mathbf{E}/\partial t$ in Maxwell equations, assuming that all velocities are much less than speed of light and the system can be described in a so-called Darwin (low-frequency) approximation Krause et al. (2007). The governing equations then become Hesse et al. (1995):
@@
where $n_{e,i}(t,\mathbf{r})$ are electron and ion densities, $\mathbf{V}_{e,i}(t,\mathbf{r})$ are averaged (hydrodynamical) electron and ion velocities, $\mathbb{P}$ is a pressure tensor of electrons, $\nabla$ is a del (nabla) vector differential operator and $(\cdot)^{T}$ means tensor transposition. We would like to underline that despite displacement current is neglected it is not a purely electrostatic approximation as we still have time-dependent electric field and inductive magnetic field. It allows to model low-frequency magneto-hydrodynamics waves and instabilities as well as Weibel-like instabilities with somewhat corrected dispersion relations. Those corrections, however, are of the order of $V_{e,i}/c$ Sarrat et al. (2016); Sladkov and Korzhimanov (2023).
The two-dimensional hybrid simulation was carried out by the AKA code Sladkov et al. (2020). This code was previously successfully tested on a model problem for analyzing the magnetic reconnection process in the perturbed Harris sheet on electron spatial scales Sladkov et al. (2021), in the problem of long-term dynamics of the Weibel instability Sladkov and Korzhimanov (2023), as well as in modeling experiments on laser ablation of plasma in an external magnetic field Bolaños et al. (2022); Burdonov et al. (2022); Zemskov et al. (2024) and in self-generated magnetic fields Sladkov et al. (2024).
@@
Initial flow velocity, $V_{0}$ $10^{6}$ cm/s* $10^{9}$ cm/s
@@
Ion inertial length, $d_{0}=c\sqrt{\varepsilon_{0}m_{i}/e^{2}N_{0}}$** 1.18 cm 1.18 cm
@@
* The flow velocity in hybrid simulations was varied. Here the typical value is given.
** Here $\varepsilon_{0}$ is an electric constant.
@@ Acknowledgements.
Acknowledgements. This research was funded by Russian Science Foundation grant number 23-12-00317. The simulations were performed on resources provided by the Joint Supercomputer Center of the Russian Academy of Sciences.

### v2

@@ I Introduction
In astrophysics, the interaction of plasma flows with magnetic structures play a major role. In particular, in the atmospheres of stars and the magnetospheres of planets, magnetic field configurations in the form of an arch, usually filled with plasma, are widespread. Such magnetized plasma arch structures have been observed on the Sun Masuda et al. (1994) and in the Earth’s magnetosphere Liu (2020). It has been shown that they play a significant role in solar flares Xia et al. (2020), as well as in various explosive processes accompanied by the reconnection of magnetic lines Yamada, Kulsrud, and Ji (2010).
@@
Despite the rich possibilities for performing various diagnostics in a laboratory experiment, many processes occurring in it have too small scales in time and space to be distinguished by modern methods. In this regard, numerical modeling is widely used to plan experiments and explain their results. As an example, we can cite a recent work on the numerical modeling of the interaction process of a plasma flow incident on a magnetic arch Wang, Zhang, and Duan (2024) in a configuration similar to the experiment on our setup Viktorov, Golubev, and Vodopyanov (2019a, b). At present, work is being carried out on the setup to study the collision process of two counter-streaming flows injected from the bases of a magnetic arch along magnetic lines. The present work is aimed at the numerical modeling of this process and the identification of the main interaction modes and the physical processes accompanying them. The main attention is paid to the emerging structures of plasma density, directly observed by its optical glow in the experiment, as well as to the processes of development of various instabilities and, in particular, the Weibel instability Weibel (1959).
@@ III Numerical methods
Numerical modeling of weakly collisional highly non-equilibrium plasma generally requires consistent consideration of kinetic effects, which can be ensured by numerical modeling of Vlasov—Maxwell system of equations Bertrand, Del Sarto, and Ghizzo (2019):
@@
In this regard, within the framework of this study, it was decided to conduct a full-scale modeling using a hybrid method, in which only ions are described kinetically, and electrons are represented as a massless neutralizing liquid. To partially account for kinetic effects, however, electrons were described in a so-called 10-moment approximation taking into account the evolution of their pressure tensor. We also neglect displacement current $\partial\mathbf{E}/\partial t$ in Maxwell equations, assuming that all velocities are much less than speed of light and the system can be described in a so-called Darwin (low-frequency) approximation Krause, Apte, and Morrison (2007). The governing equations then become Hesse, Winske, and Kuznetsova (1995):
@@
where $n_{e,i}(t,\mathbf{r})$ are electron and ion densities, $\mathbf{V}_{e,i}(t,\mathbf{r})$ are averaged (hydrodynamical) electron and ion velocities, $\mathbb{P}$ is a pressure tensor of electrons, $\nabla$ is a del (nabla) vector differential operator and $(\cdot)^{T}$ means tensor transposition. We would like to underline that despite displacement current is neglected it is not a purely electrostatic approximation as we still have time-dependent electric field and inductive magnetic field. It allows to model low-frequency magneto-hydrodynamics waves and instabilities as well as Weibel-like instabilities with somewhat corrected dispersion relations. Those corrections, however, are of the order of $V_{e,i}/c$ Sarrat, Del Sarto, and Ghizzo (2016); Sladkov and Korzhimanov (2023).
The two-dimensional hybrid simulation was carried out by the AKA code Sladkov, Smets, and Korzhimanov (2020). This code was previously successfully tested on a model problem for analyzing the magnetic reconnection process in the perturbed Harris sheet on electron spatial scales Sladkov et al. (2021), in the problem of long-term dynamics of the Weibel instability Sladkov and Korzhimanov (2023), as well as in modeling experiments on laser ablation of plasma in an external magnetic field Bolaños et al. (2022); Burdonov et al. (2022); Zemskov et al. (2024) and in self-generated magnetic fields Sladkov et al. (2024).
@@
Initial flow velocity, $V_{0}$ $10^{6}$ cm/s [^1] The flow velocity in hybrid simulations was varied. Here the typical value is given. $10^{9}$ cm/s
@@
Ion inertial length, $d_{0}=c\sqrt{\varepsilon_{0}m_{i}/e^{2}N_{0}}$ [^2] Here $\varepsilon_{0}$ is an electric constant. 1.18 cm 1.18 cm
@@ Acknowledgements.
This research was funded by Russian Science Foundation grant number 23-12-00317. The simulations were performed on resources provided by the Joint Supercomputer Center of the Russian Academy of Sciences.